OpenStudy (idealist10):

Find the orthogonal trajectories of xye^(x^2)=C.

OpenStudy (anonymous):

0

OpenStudy (idealist10):

\[x(2xye ^{x^2}+\frac{ dy }{ dx }e ^{x^2})+ye ^{x^2}=0\]

OpenStudy (anonymous):

solve it!

OpenStudy (idealist10):

\[2x^2ye ^{x^2}+\frac{ dy }{ dx }xe ^{x^2}+ye ^{x^2}=0\]

OpenStudy (idealist10):

\[\frac{ dy }{ dx }xe ^{x^2}=-2x^2ye ^{x^2}-ye ^{x^2}\]

OpenStudy (anonymous):

UH

OpenStudy (idealist10):

\[\frac{ dy }{ dx }=-2xy-\frac{ y }{ x }=\frac{ -2x^2y-y }{ x }\]

OpenStudy (idealist10):

\[\frac{ dy }{ dx }=\frac{ -y(2x^2+1) }{ x }\]

OpenStudy (idealist10):

negative reciprocal of dy/dx is...\[\frac{ dy }{ dx }=\frac{ x }{ y(2x^2+1) }\]

OpenStudy (idealist10):

\[y dy=\frac{ x }{ 2x^2+1 }dx\]

OpenStudy (idealist10):

u=2x^2+1 du=4x dx (1/4)du=x dx \[\frac{ 1 }{ 4 }\ln \left| 2x^2+1 \right|+C\]

OpenStudy (idealist10):

\[\frac{ y^2 }{ 2 }=\frac{ 1 }{ 4 }\ln \left| 2x^2+1 \right|+C\]

OpenStudy (idealist10):

\[y^2=\frac{ 1 }{ 2 }\ln \left| 2x^2+1 \right|+C\]

OpenStudy (idealist10):

But the answer in the book is \[y^2=-\frac{ 1 }{ 2 }\ln(1+2x^2)+k\]

OpenStudy (idealist10):

And I don't know which answer is right and why the book has the negative sign in front of 1/2.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

@satellite73

OpenStudy (anonymous):

It's probably a typo. Your work is correct.

OpenStudy (idealist10):

Thank you for verifying my work.